# Graph Permutation Entropy: Extensions to the Continuous Case, A step towards Ordinal Deep Learning, and More

### v1

@@ Abstract
Index Terms:
Graph signals, Continuous permutation entropy, Ordinal activation function, Graph ordinal patterns, Ordinal contrasts, multimodal MRI.
@@ I Introduction
Here, we extend the idea behind this methodology to the computation of a continuous version of $PE_{G}$. Our approach adapts a more natural method of describing continuous ordinal patterns and uses their mutual exclusivity as a way to compute a joint probability distribution. In particular, we introduce a permutation activation function. This takes any continuous, differentiable and strictly monotonically increasing function that maps embedding vectors to a continuous domain while maintaining the ordinal structure of the patterns. This allows us to exploit certain amplitude-related features in windows of a given length depending on the choice of activation function. We show how this is an improvement on the discrete case in detecting non-linear spikes in signals in real-life and synthetic data.
@@
I-A Contributions
In sum, the contributions of this study are:
  • Introduction of a Continuous Version of Graph Permutation Entropy.
  • The development of the permutation activation function, a novel step towards ordinal deep learning.
  • Introduction of ordinal pattern contrasts of length 3 formally to the graph domain.
  • Development of continuous graph permutation pattern contrasts of length 3.
  • Introduction of 2 new pattern contrasts for the ordinal analysis of images in the graph domain.
  • Validation of developed methodologies by analyzing non-linear dynamics in synthetic and real life data.
@@ II-E Continuous (Graph) Permutation Entropy and Ordinal Contrasts
We first define our permutation activation function. This is a function that adds non-linearity to the embedding vector in Graph Permutation Entropy. This corresponds to the traditional deep learning activation function. Note this can be applied to traditional Permutation Entropy analogously.
@@
    Apply the permutation activation function ($\rho_{\text{activate}}$).
    $\textbf{a}^{m,L}_{i}$ $=\rho_{\text{act}}(y^{m,L}_{i}).$
@@
    For each permutation of length $m$, we store the corresponding continuous permutation activated vector $a$ in the respective pattern class. The new embedding vector is now:
@@
As this property translates trivially to $PE_{G}$. It is implied that the application of our permutation activation function on windows of length $m$ preserves the Discrete $PE_{G}$ of the signal in all cases.
@@
As a possible pre-processing step, $z$-score normalization can be used over the whole graph signal before constructing the embedding vector to account for the case where signal values are extremely small or extremely large. This can result in every value in the signal being mapped to the same value due to the structure of certain permutation activation functions. Using the hyperbolic tangent as an example with the range $[-1,1]$, very large values would be mapped to 1 and vice-versa for very small values. Now everything will be mapped to the class accounting for ties making the method impractical. This is prominent in some signal types, such as fMRI signals where signal values are typically very large. Interestingly, this can be seen as a direct ordinal deep learning translation of the vanishing/exploding gradient problem. (Optional) Normalization over the neighbourhood topology ensures that while the range of the signal is still technically infinite, extreme values are very unlikely to occur. The normalization is typically done over the whole signal x at the start so relative amplitudes are maintained in the windows. Once again, this is similar to deep learning methods where normalization is sometimes used as a pre-processing step [9]. We can use any type of normalization such as min-max normalization as well depending on the data at hand. As in the discrete case, each node has a continuous pattern maintaining the granular analysis efficiency. Ties are very rare due to how the embedding vector is constructed by taking neighbourhood averages, the continuous permutation activation function being strictly monotonically increasing further reduces the likelihood of ties, in-fact in our experiments we did not encounter any ties.
The use of a statistic ($T$) allows us to extract information from the activated window that we deem relevant. For example, we may use the variance when we think the variation in amplitude may have discriminating information, or the mean when the entire window may be informative. Using the median as $T$ for a non-negative monotonically increasing permutation activation function returns the intermediate value method used by Zanin [10], while using the rank function (known as tiedrank in MATLAB) as our permutation activation function with the identity function (maps all vectors to a scalar value of 1), we obtain classical $PE_{G}$. Given that these operations are consistent over all windows, the following step to compute the relative weighted probabilities and thus the probability distribution for input into Shannon Entropy is a valid operation. In essence, we are projecting all of the embedding vectors from the real line to the same permutation activation function and comparing their relative behaviour on this continuous function.
@@
Consider an embedding vector where $m=3$: $(0.5,0.1,1.7)$. Now let us choose a permutation activation function. For example, the GELU (Gaussian Error Linear Unit) [16] function, a commonly used activation function in many deep learning use cases, is defined as:
@@
with $\mathrm{erf}$ being the traditional Gaussian error function. Applying this function gives us $(0.3457,0.0540,1.6242)$. Now we choose the mean to be our statistic ($T$) and we compute $\left|T(a^{m,L}_{i})|\right|$. This gives us 0.6747 as our $*_{i,p}$ value for that node. We must take the absolute value as having negative weights in our probability computation will result in incorrect results. This can be considered as a weighted ‘count’ as in discrete $PE_{G}$ we would have the value just being 1, we have 0.2282 here, determined by the non-linearity of the permutation activation function and the corresponding statistic. These are parameters that can be fine-tuned based on a priori information about the signal under consideration. Note how throughout the computation the order (213) of the embedding vector is maintained.
@@
Observe that now the relative pattern probability distributions are being changed and mapped in a non-linear way while still maintaining the nature of the ordinal patterns. The higher amplitude pattern windows are being mapped to higher values at a rate determined by the permutation activation function. Another way to interpret this is that instead of measuring the proportion each ordinal pattern makes up of the total number of patterns we are measuring how much of the total power of the (graph) signal each ordinal pattern is contributing to.
Theoretically, we can now manipulate the parameters of this algorithm to ‘train’ the model on a given signal or learn the best permutation activation function specific to the signal. For example, if we know that most values in the signal fluctuate around a certain range but relevant signal spikes are also not too different from this general background amplitude, we could use an exponentially increasing function to emphasize larger values or spikes in the neighbourhood as illustrated in Figure 6. We can also now optimize the gradient of the permutation activation function as an iterative training step similar to back-propagation. Furthermore, not only are amplitudes considered within windows but relative to the whole signal as well, addressing a long-running drawback of permutation entropy. The permutation activation function can also be optimized to reduce the influence of noise by mapping to a function with a very slowly increasing gradient maintaining a stable value. This can clearly be extended to uni-variate and multi-variate time-series signals.
Figure 2: Example of spike emphasis when using exp permutation activation function
@@
This is a straightforward extension as the probabilities can just be considered permutation activation function weighted relative pattern frequencies compared to the relative pattern frequencies in the discrete case.
@@ III Results
In Figure 4 we notice that the exponential function is the one providing results most similar to the discrete version of $PE_{G}$. The hyperbolic tangent and GELU function seems to deviate more from the discrete case and can detect different islands of stability more clearly. This example illustrates how tuning the permutation activation function can highlight various signal features of interest. While such tuning is beneficial, particularly in the context of deep learning, we must be cautious to avoid data dredging or overfitting. Importantly, even with adjustments to the activation function, the overall trends remain apparent in most cases.
@@
Exploring this further, we assessed how much the probability distribution of $CPE_{G}$ deviated from $PE_{G}$ at different values of $r$ in the logistic map for different permutation activation functions. For this, we use the Kullback-Leibler Divergence [24] as our measure of deviation. The Kullback-Leibler Divergence (KL divergence) between two probability distributions $P$ and $Q$ is defined as:
@@
Inspecting Figure 5, all 3 permutation activation functions remain stable for chaotic behaviour, i.e., their distributions are relatively similar to the discrete case. We note that functions with faster increasing gradients seem to be more stable relative to Discrete $PE_{G}$, while more slowly increasing gradients results in more deviation from the discrete case with large deviations particularly in the islands of stability. In fact, the computation of the KL divergence between such functions and discrete $PE_{G}$ can actually detect islands of stability within chaotic behaviour. This implies that while the patterns that dominate the total frequency of patterns are similar to the patterns that contribute most to the overall power of the signal at islands of stability this is no longer the case.
It is useful to observe that in both the continuous $PE_{G}$ plot and the KL-Divergence plot that the global behaviour is very similar with subtle changes due to choice of permutation activation function (the statistic is kept constant as the mean) with most deviation in behavior occurring at the islands of stability. This is as expected and highlights how using different activation functions alters the behaviour at change-points allowing us to emphasize those that we deem relevant.
@@ III-B Detecting non-linear spikes using Continuous Permutation Entropy
We use as our permutation activation functions the exponential function $e$, and the sigmoid function:
@@
Using an undirected path graph to map the time-series as a graph signal, we computed the $PE_{G}$ and contrast values for both permutation activation functions and also for the discrete case for each of the 100 realizations of the signals with and without the Gaussian spike. Essentially we have two groups of 100 values (for each contrast and $(C)PE_{G}$) one with the spike and one without. We then performed the Wilcoxon rank-sum test for significance between the two groups.
@@ III-C Real-world example: Analysis of weather data
Table V shows our results. The exponential permutation activation function performs the best in terms of $PE_{G}$, picking up the more rapid increases in temperature very well. In fact, all of its contrasts, aside from $\gamma$, picked up the difference. In the discrete case, only the $\rho$ contrast detected the spike, while in the hyperbolic tangent case the $\delta$ and $\beta$ contrasts picked up the spike. Additionally, it seems the $\beta$ contrasts performance improves with permutation activation functions with larger gradients. Conversely, the $\delta$ contrast appears to detect the change regardless of the underlying permutation activation function.
@@ III-D Heartbeat time-series
Then, we compute the $CPE_{G}$ on both the surrogate and normal graph signals form the heartbeat data. We perform the non-parametric Wilcox on Rank-sum test for significance between each patients signal and its surrogate. We compute the $CPE_{G}$ and the continuous contrasts with the exponential permutation activation function for all 6 windows over all participants and evaluate if they can distinguish between the respective participant signal and its surrogate. We repeat this for the discrete case.
@@ IV Discussion
We developed a continuous version of Graph Permutation Entropy with a novel permutation activation function. This strictly monotonically increasing function can emphasize certain features of a given signal (spikes) or suppress undesired ones (noise) while maintaining the ordinal structure of the embedding vector and keeping the overall number of each pattern permutation the same. We showed there are many ways we can optimize the activation permutation function and the statistic we use to extract window information. This creates a link between $CPE_{G}$ and deep learning by being able to optimize parameters to get a desired result. Note that this method, similar to Zanin [10] creates another graph signal with $N$ nodes (or an $N$ length signal in the uni-variate case). This can be considered a convolution filtering operation. Thus, there is now a way to learn the optimal ‘filter’ by exploring an infinite search space of monotonically increasing functions. In terms of further applications to machine learning or graph neural networks [40], this is promising.
@@ V Conclusions and Future Work
$CPE_{G}$, the continuous counterpart of $PE_{G}$, marks a significant advancement in ordinal deep learning, extending the groundwork laid by Zanin. We introduce the permutation activation function, enabling exploration within the infinite search space of strictly monotonically increasing functions. While strictly monotonically decreasing functions could maintain global consistency with $PE_{G}$ as per the in-variance theorem, their use may compromise interpretability in ordinal contrasts and granular node-level analyses due to the reversal in order. This undermines $CPE_{G}$’s hallmark feature of being able to maintain the distinct ordinal patterns of each node from classical discrete $PE_{G}$ while adding extra continuous information.
@@
It is vital to note that this work is not merely an adaptation of previous amplitude-aware extensions of Permutation Entropy, but a novel general framework that uses the projection of embedding vectors onto a continuous permutation activation function to get more fine-grained detail on the signal’s non-linear dynamics while maintaining the simplicity of the ordinal patterns in traditional PE.

### v2

@@ Abstract
Understanding complex data requires advanced methods, especially for signals from irregular domains like graphs. This research introduces a continuous version of Permutation Entropy, adapting it for graph signals to enable nonlinear analysis. By developing an ordinal activation function and extending ordinal contrasts to graphs, the study advances the emerging field of Ordinal Deep Learning. Experiments on synthetic and real-life data across multiple disciplines demonstrate the effectiveness of these methods, promising enhanced data analysis and interpretability.
@@ I Introduction
Here, we extend the idea behind this methodology to the computation of a continuous version of $PE_{G}$. Our approach adapts a more natural method of describing continuous ordinal patterns and uses their mutual exclusivity as a way to compute a joint probability distribution. In particular, we introduce a ordinal activation function. This is any continuous and strictly monotonically increasing function that maps embedding vectors to a continuous domain while maintaining the ordinal structure of the patterns. This allows us to exploit certain amplitude-related features in windows of a given length depending on the choice of activation function. We show this is an improvement on the discrete case in real-life and synthetic data.
@@ II-E Continuous (Graph) Permutation Entropy and Ordinal Contrasts
We first define our ordinal activation function. This is a function that adds non-linearity to the embedding vector in Graph Permutation Entropy. This corresponds to the traditional deep learning activation function. Note this can be applied to traditional Permutation Entropy analogously.
@@
    Apply the ordinal activation function ($\rho_{\text{act}}$).
    $\textbf{a}^{m,L}_{i}$ $=\rho_{\text{act}}(\textbf{y}^{m,L}_{i}).$
@@
    For each permutation of length $m$, we store the corresponding continuous ordinal activated vector $a$ in the respective pattern class. The new embedding vector is now:
@@
As this property translates trivially to $PE_{G}$. It is implied that the application of our ordinal activation function on windows of length $m$ preserves the Discrete $PE_{G}$ of the signal in all cases.
@@
As a possible pre-processing step, $z$-score normalization can be used over the whole graph signal before constructing the embedding vector to account for the case where signal values are extremely small or extremely large. This can result in every value in the signal being mapped to the same value due to the structure of certain ordinal activation functions. Using the hyperbolic tangent as an example with the range $[-1,1]$, very large values would be mapped to 1 and vice-versa for very small values. Now everything will be mapped to the class accounting for ties making the method impractical. This is prominent in some signal types, such as fMRI signals where signal values are typically very large. Interestingly, this can be seen as a direct ordinal deep learning translation of the vanishing/exploding gradient problem. (Optional) Normalization over the neighbourhood topology ensures that while the range of the signal is still technically infinite, extreme values are very unlikely to occur. The normalization is typically done over the whole signal x at the start so relative amplitudes are maintained in the windows. Once again, this is similar to deep learning methods where normalization is sometimes used as a pre-processing step [9]. We can use any type of normalization such as min-max normalization as well depending on the data at hand. As in the discrete case, each node has a continuous pattern maintaining the granular analysis efficiency. Ties are very rare due to how the embedding vector is constructed by taking neighbourhood averages, the continuous ordinal activation function being strictly monotonically increasing further reduces the likelihood of ties, in-fact in our experiments we did not encounter any ties.
The use of a statistic ($T$) allows us to extract information from the activated window that we deem relevant. For example, we may use the variance when we think the variation in amplitude may have discriminating information, or the mean when the entire window may be informative. Using the median as $T$ for a non-negative monotonically increasing ordinal activation function returns the intermediate value method used by Zanin [10], while using the rank function (known as tiedrank in MATLAB) as our ordinal activation function with the identity function (maps all vectors to a scalar value of 1), we obtain classical $PE_{G}$. Given that these operations are consistent over all windows, the following step to compute the relative weighted probabilities and thus the probability distribution for input into Shannon Entropy is a valid operation. In essence, we are projecting all of the embedding vectors from the real line to the same ordinal activation function and comparing their relative behaviour on this continuous function.
@@
Consider an embedding vector where $m=3$: $(0.5,0.1,1.7)$. Now let us choose a ordinal activation function. For example, the GELU (Gaussian Error Linear Unit) [16] function, a commonly used activation function in many deep learning use cases, is defined as:
@@
with $\mathrm{erf}$ being the traditional Gaussian error function. Applying this function gives us $(0.3457,0.0540,1.6242)$. Now we choose the mean to be our statistic ($T$) and we compute $\left|T(a^{m,L}_{i})|\right|$. This gives us 0.6747 as our $*_{i,p}$ value for that node. We must take the absolute value as having negative weights in our probability computation will result in incorrect results. This can be considered as a weighted ‘count’ as in discrete $PE_{G}$ we would have the value just being 1, we have 0.2282 here, determined by the non-linearity of the ordinal activation function and the corresponding statistic. These are parameters that can be fine-tuned based on a priori information about the signal under consideration. Note how throughout the computation the order (213) of the embedding vector is maintained.
@@
Observe that now the relative pattern probability distributions are being changed and mapped in a non-linear way while still maintaining the nature of the ordinal patterns. The higher amplitude pattern windows are being mapped to higher values at a rate determined by the ordinal activation function. Another way to interpret this is that instead of measuring the proportion each ordinal pattern makes up of the total number of patterns we are measuring how much of the total power of the (graph) signal each ordinal pattern is contributing to.
Theoretically, we can now manipulate the parameters of this algorithm to ‘train’ the model on a given signal or learn the best ordinal activation function specific to the signal. For example, if we know that most values in the signal fluctuate around a certain range but relevant signal spikes are also not too different from this general background amplitude, we could use an exponentially increasing function to emphasize larger values or spikes in the neighbourhood as illustrated in Figure 6. We can also now optimize the gradient of the ordinal activation function as an iterative training step similar to back-propagation. Furthermore, not only are amplitudes considered within windows but relative to the whole signal as well, addressing a long-running drawback of permutation entropy. The ordinal activation function can also be optimized to reduce the influence of noise by mapping to a function with a very slowly increasing gradient maintaining a stable value. This can clearly be extended to uni-variate and multi-variate time-series signals.
Figure 2: Example of spike emphasis when using exp ordinal activation function
@@
This is a straightforward extension as the probabilities can just be considered ordinal activation function weighted relative pattern frequencies compared to the relative pattern frequencies in the discrete case.
@@ III Results
In Figure 4 we notice that the exponential function is the one providing results most similar to the discrete version of $PE_{G}$. The hyperbolic tangent and GELU function seems to deviate more from the discrete case and can detect different islands of stability more clearly. This example illustrates how tuning the ordinal activation function can highlight various signal features of interest. While such tuning is beneficial, particularly in the context of deep learning, we must be cautious to avoid data dredging or overfitting. Importantly, even with adjustments to the activation function, the overall trends remain apparent in most cases.
@@
Exploring this further, we assessed how much the probability distribution of $CPE_{G}$ deviated from $PE_{G}$ at different values of $r$ in the logistic map for different ordinal activation functions. For this, we use the Kullback-Leibler Divergence [24] as our measure of deviation. The Kullback-Leibler Divergence (KL divergence) between two probability distributions $P$ and $Q$ is defined as:
@@
Inspecting Figure 5, all 3 ordinal activation functions remain stable for chaotic behaviour, i.e., their distributions are relatively similar to the discrete case. We note that functions with faster increasing gradients seem to be more stable relative to Discrete $PE_{G}$, while more slowly increasing gradients results in more deviation from the discrete case with large deviations particularly in the islands of stability. In fact, the computation of the KL divergence between such functions and discrete $PE_{G}$ can actually detect islands of stability within chaotic behaviour. This implies that while the patterns that dominate the total frequency of patterns are similar to the patterns that contribute most to the overall power of the signal at islands of stability this is no longer the case.
It is useful to observe that in both the continuous $PE_{G}$ plot and the KL-Divergence plot that the global behaviour is very similar with subtle changes due to choice of ordinal activation function (the statistic is kept constant as the mean) with most deviation in behavior occurring at the islands of stability. This is as expected and highlights how using different activation functions alters the behaviour at change-points allowing us to emphasize those that we deem relevant.
@@ III-B Detecting non-linear spikes using Continuous Permutation Entropy
We use as our ordinal activation functions the exponential functionand the sigmoid function:
@@
Using an undirected path graph to map the time-series as a graph signal, we computed the $PE_{G}$ and contrast values for both ordinal activation functions and also for the discrete case for each of the 100 realizations of the signals with and without the Gaussian spike. Essentially we have two groups of 100 values (for each contrast and $(C)PE_{G}$) one with the spike and one without. We then performed the Wilcoxon rank-sum test for significance between the two groups.
@@ III-C Real-world example: Analysis of weather data
Table V shows our results. The exponential ordinal activation function performs the best in terms of $PE_{G}$, picking up the more rapid increases in temperature very well. In fact, all of its contrasts, aside from $\gamma$, picked up the difference. In the discrete case, only the $\rho$ contrast detected the spike, while in the hyperbolic tangent case the $\delta$ and $\beta$ contrasts picked up the spike. Additionally, it seems the $\beta$ contrasts performance improves with ordinal activation functions with larger gradients. Conversely, the $\delta$ contrast appears to detect the change regardless of the underlying ordinal activation function.
@@ III-D Heartbeat time-series
Then, we compute the $CPE_{G}$ on both the surrogate and normal graph signals form the heartbeat data. We perform the non-parametric Wilcox on Rank-sum test for significance between each patients signal and its surrogate. We compute the $CPE_{G}$ and the continuous contrasts with the exponential ordinal activation function for all 6 windows over all participants and evaluate if they can distinguish between the respective participant signal and its surrogate. We repeat this for the discrete case.
@@ IV Discussion
We developed a continuous version of Graph Permutation Entropy with a novel ordinal activation function. This strictly monotonically increasing function can emphasize certain features of a given signal (spikes) or suppress undesired ones (noise) while maintaining the ordinal structure of the embedding vector and keeping the overall number of each pattern permutation the same. We showed there are many ways we can optimize the activation permutation function and the statistic we use to extract window information. This creates a link between $CPE_{G}$ and deep learning by being able to optimize parameters to get a desired result. Note that this method, similar to Zanin [10] creates another graph signal with $N$ nodes (or an $N$ length signal in the uni-variate case). This can be considered a convolution filtering operation. Thus, there is now a way to learn the optimal ‘filter’ by exploring an infinite search space of monotonically increasing functions. In terms of further applications to machine learning or graph neural networks [40], this is promising.
@@ V Conclusions and Future Work
$CPE_{G}$, the continuous counterpart of $PE_{G}$, marks a significant advancement in ordinal deep learning, extending the groundwork laid by Zanin. We introduce the ordinal activation function, enabling exploration within the infinite search space of strictly monotonically increasing functions. While strictly monotonically decreasing functions could maintain global consistency with $PE_{G}$ as per the in-variance theorem, their use may compromise interpretability in ordinal contrasts and granular node-level analyses due to the reversal in order. This undermines $CPE_{G}$’s hallmark feature of being able to maintain the distinct ordinal patterns of each node from classical discrete $PE_{G}$ while adding extra continuous information.
@@
It is vital to note that this work is not merely an adaptation of previous amplitude-aware extensions of Permutation Entropy, but a novel general framework that uses the projection of embedding vectors onto a continuous ordinal activation function to get more fine-grained detail on the signal’s non-linear dynamics while maintaining the simplicity of the ordinal patterns in traditional PE.